XY=14 and x+y=9 then x-y=

Answer by anirudhreddy(2)

XY=14 and x+y=9 then x-y=
to solve the above problem we need to learn basic formulas like 1)(a+b)^2= a^2+b^2+2ab. and 2)(a-b)^2=a^2+b^2-2ab.
now apply the above formulas to given data by squaring both sides for x+y=9
==> (x+y)^2=9^2 ==> x^2+y^2+2xy=81, now assign given assign values to xy
==> x^2+y^2+2(14)=81 ==> x^2+y^2= 81-28. ==> x^2+y^2= 53.
now apply this x^2+y^2 values to formula 2. (x-y)^2=x^2+y^2-2xy.
==> (x-y)^2= 53-28 ==> (x-y)^2= 25. ==> now taking sq.root on both sides and we get the x-y values as x-y=5.